OpenStudy (anonymous):

Using substitution formula to evaluate the integral ^pi/4[0 (1-cos2t)sin2t dt?

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ \pi }{ 4 }} (1-\cos2t)\sin2t dt\]

hartnn (hartnn):

so did you try substituting u = 1-cos 2t ?

OpenStudy (anonymous):

so far I have u=1-cos2t) and du=sin2t dt Can someone tell me if I have this substation done correctly? From there I have t=0; 1-cos2(0)=0 t(pi/4)=1-cos2(pi/4)= and I seem to be stuck right here if my above equation is correct.

hartnn (hartnn):

correct substitution :)

hartnn (hartnn):

cos 2(pi/4) = cos pi/2 = ... ?

OpenStudy (anonymous):

Oh I see I did not cross cancel the two from pi/4 . Thank you. Now it should be \[\int\limits_{0}^{\pi/4}udu=\int\limits_{0}^{\pi/4} u(\sin2t) \frac{ du }{ 2\sin2t }\] now the (sin2t) crosses out with the fraction of du/2 sin2t leaving 1/2. Now I did get done talking to my teacher about this problem and she said the problem should look like this \[\int\limits_{0}^{\pi/4} \frac{ 1 }{ 2}*\frac{ u^2 }{ 2 }\] can someone explain where the u^2/2 came from?

hartnn (hartnn):

the problem should look like \(\Large \int \limits _0^1 (u/2) du\)

hartnn (hartnn):

while doing substitution, we change the limits also!

hartnn (hartnn):

so, u = 1-cos 2t 0--->0 pi/4 --->1 du/2 = sin 2t dt clear till here ? substitute all these in you integral to form a new easier integral

hartnn (hartnn):